OpenStudy (anonymous):

i give medals if you help One person can do a certain job in ten minutes, and another person can do the same job in fifteen minutes. How many minutes will they take to do the job together? 4 min 5 min 6 min

OpenStudy (raden):

x / +

OpenStudy (kewlgeek555):

A can do the job in 10 minutes or, In 1 minute, A can do 1/10 of the job. B can do the job in 15 minutes or, In 1 minute, B can do 1/15 of the job. Together, in 1 minute they can do 1/10+1/15 = 1/6 of the job. So, it would take 6/1 = 6 minutes to finish the job if they both work together.

OpenStudy (kewlgeek555):

Where A is the one person and B is the "Another Person".

OpenStudy (raden):

(10 * 15) / (10+15) = 6

OpenStudy (anonymous):

thanks

OpenStudy (raden):

yw